Systems Staff Meeting Notes 6/19/00

Systems Staff Meeting Notes
June 19, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Ying has accepted the DLC programmer position starting today. Now we'll have to hire her replacement and perhaps even another programmer position.
  2. Systems security over the weekend - Rich noticed some random lights left on in the hallway and the door by the stairwell unlocked. Hilton notes that his computer was turned on. Everett was here over the weekend. Will came in about midnight on Friday to reboot NT9. Staff should check their areas to see if anything is missing, and keep an eye out this week.
  3. Halon fire suppression people will be here this evening to inspect our system in the tape library.
  4. During Bill's leave of absence in the fall, Chris Leopold (IFAS) will be our consultant on call for about 10 hours per week.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. We've received nine more black PDA's. We're missing eight blue ones. Martha has given Bill a distribution list; will try to dispense them today. Bill has experimented & determined that you can use an Excel spreadsheet with them. They could also be handy for recording network configuration information and are especially useful for maintaining a personal calendar. We can probably expect to get some for the Systems department.
  2. Switch outage on 223 last Friday - North Tower where it feeds from the building POP. The switch dumped its memory; Will replaced it with another switch from the vault. The old switch will be returned once it is properly configured.
Networking
  1. NT9 - is currently up & running fine. Will put Hot Fix on it last Friday - none of the printers would work. He took Hot Fix off and now the printers are fine. Will is planning to contact Microsoft about the problem. The hot fixes are not cumulative; they've had less regression testing. The service packs are cumulative, but aren't even getting released quarterly.
  2. West wiring - the job is finished, Digital is out of the building. They are currently working in MSL. There were some complaints about the lack of warning regarding the installation process over there. Any time we know that they'll be starting a project, we need to alert the affected departments in advance.
  3. We have final estimates for the DLC & Resource Services changes. Estimates are pending on our work in West, we have to check the status of the work that Digital just finished up.
  4. We had a complaint in 110A Maps regarding the noise generated by the switch. We will rectify the situation since we have the money to do so - now rewiring projects will pull everything punched down back to the closets.
  5. We'll be giving some of the old, pink wire to Digital, saving a few boxes for us. The wire mold we have can go to Facilities.
  6. Wireless - Will received an email saying that the conduit for the electrical has been run. Should be available in the center ceiling of the third floor stacks. Should have pulled two, Will needs to check in case he only pulled one.
  7. Novell CD conversion - will take place next week.
  8. We'll need to relocate the servers to room 229 West.
Software Projects Status
  1. Q&A replacement - Cathy Mook has requested some slight modifications. They are still testing the database.
  2. Registry Migration - should be finished this afternoon.
  3. Library Statistics - some changes have been made to accommodate the user's request to not have to manually refresh. It now does it automatically at the server level.
  4. LASSIE is finished - it needed a sorting option. They are very pleased with it now.
  5. IR database - cube testing has not yet been done. Time must be allotted for this, even if it means putting other projects on the back burner temporarily.
DP Coordinator
  1. Student will be out on intercession break next week, June 26-30.
  2. Becky will be gone after this Friday; Kelly is next at the end of Summer B. Will be seeking replacements.
  3. Lindsay's last day is this Thursday. Her replacement, Andy, seems to be working out.
Calendar
  1. Winston is out all this week.
  2. Hilton will be out July 3.
  3. Independence Day holiday is Tuesday, July 4.
  4. Bobbie will be out July 10-21.
  5. Will out July 24-28.
  6. ALA is July 7-11.
  7. Rich will be out July 5-25.


Back to the Systems Homepage

Last updated June 26, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu